OpenStudy (anonymous):

Evaluate the indefinite integrals... (cot5x+csc5x)dx

OpenStudy (australopithecus):

cos(5x)/sin(5x) + 1/sin(5x)dx = cos(5x) + 1/sin(5x)dx u = cos(5x) + 1 du = -5sin(5x)dx -du/5 = sin(5x)dx -1/5 integral of u dx = -(ln|cos(5x) + 1| + c)/5 waiting for someone to correct me

OpenStudy (australopithecus):

Looks Right to me actually

OpenStudy (australopithecus):

the sin(5x) is canceled out

OpenStudy (australopithecus):

in the denominator also https://www.wolframalpha.com/input/?i=%28cot5%2Bcsc5%29+%3D+%28cos%285%29+%2B+1%29%2Fsin%285%29 proof that your initial problem can be simplified to what I simplified it to

OpenStudy (australopithecus):

wait never mind I made a mistake

OpenStudy (anonymous):

you almost correct. actually i already have a correct answer for that problem. My problem is I dont know the solution. :)

OpenStudy (anonymous):

this is the correct answer = 1/5ln(1-cos5x)+C will you show me the solutions of it? :)

OpenStudy (australopithecus):

cos(5x)/sin(5x) + 1/sin(5x)dx = cos(5x) + 1/sin(5x)dx u = cos(5x) + 1 du = -5sin(5x)dx -du/5sin(5x) = dx So we have -1/5 integral of udu/sin^(2)(5x) write in terms of u so arccos(u-1)/ = 5x so -1/5 integral of udu/(1-cos^(2)(arcos(u-1)) = -1/5 integral of udu/(1-(u-1)^(2)) = -1/5 integral of udu/1-(u^(2)-2u+1) = -1/5 integral of udu/1-u^(2)+2u-1 = -1/5 integral of udu/(-u^(2)+2u) = -1/5 integral of udu/1-(u^(2)-2u+1) = -1/5 integral of udu/(-u^(2)+2u) = -1/5 integral of du/(-u+2) Now another round of substitution I get s = -u+2 ds = -1du -ds = du thus 1/5 integral of ds/s = (1/5)ln|s| + c sub back in (1/5)ln|-u+2| + c sub back in (1/5)ln|-cos(5x) - 1 + 2| + c = (1/5)ln|-cos(5x) + 1| + c

OpenStudy (australopithecus):

there is your answer it is a very dirty integral if you ask me lol

OpenStudy (anonymous):

nice one @Australopithecus. Thanks a lot for your solutions. :)

OpenStudy (australopithecus):

if you have any questions about my solution feel free to ask It is kind of messy